OpenStudy (anonymous):

Could someone please help me?!?!! Squared rooted28+7+3x2divided by 3

OpenStudy (anonymous):

\[\sqrt{28+7}+3^{2}\div3\]

OpenStudy (anonymous):

5.9+9/3

OpenStudy (anonymous):

9/3=3 5.9+3=8.9 Answer=8.9

OpenStudy (anonymous):

Are you sure thats right?

OpenStudy (anonymous):

\[\sqrt{28+7}+3^2÷3\] √35+9÷3 3+ √35

OpenStudy (anonymous):

depending on if you need a decimal or the radical

OpenStudy (anonymous):

well the square root of 35 is 5.9160797831

OpenStudy (anonymous):

so i just did 5.9

OpenStudy (anonymous):

Radical

OpenStudy (anonymous):

well normally you would simplify the radical like √8 would be 2√2 but because √35 can't be simplified you would leave it as is

OpenStudy (anonymous):

yes that is correct cha1234

OpenStudy (anonymous):

@Cha1234 so then your answer is right?

OpenStudy (anonymous):

it depends on what answer you need

OpenStudy (anonymous):

3+the square root of 35 is correct

OpenStudy (anonymous):

3+ √35

OpenStudy (anonymous):

I need the radical answer

OpenStudy (anonymous):

well 3+ √35 is your correct answer :)

OpenStudy (anonymous):

I dont know...:/